# Bethe equations for the critical three-state Potts spin chain with toroidal boundary conditions

### v1

@@ 4 The Bethe equation for $T^{(c)}(x)$
where now the number of sine terms is the same as that present in the general form (33) for the transfer matrix elements. We arrive at this conclusion by inspecting the asymptotic limit of the eigenvalues of $T^{(c)}(x)$ at the points $\pm i\infty$ for small lattice sizes.
As before, the Bethe equation for the variables $\xi_{k}$ is uncovered with the help of a functional relation for the eigenvalue involving distinct spectral parameters. This equation is derived from a matrix identity for the product of transfer matrices whose main structure is similar to that presented in previous section, see Eq.(37). More specifically, we find that $T^{(c)}(x)$ satisfies the following matrix identity,
$T^{(c)}(x-\frac{\pi}{3})T^{(c)}(x-\frac{\pi}{6})T^{(c)}(x)$ $=$ $T^{(c)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T^{(c)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T^{(c)}(x)$ (48)
- $\big[f_{3}(x)\big]^{L}T^{(c)}(x+\frac{\pi}{3})\bigg\}$
where the main difference with Eq.(37) is the presence of a minus sign on the term proportional to the amplitude $f_{3}(x)$.
@@ 4.1 Bethe roots for L=2
We now proceed to investigate the Bethe equation parameterization we have found for the spectrum of the Hamiltonian $H^{(c)}$ with $L=2$. This Hamiltonian has a $Z(2)$ invariance and the Hilbert space splits into two sectors labeled by the eigenvalues of the charge ${\cal{O}}_{Z(2)}$. If we denote these eigenvalues by the symbol $\nu_{c}$ we find that there are $(3^{L}+1)/2$ states with $\nu_{c}=+1$ and $(3^{L}-1)/2$ states with $\nu_{c}=-1$. We obtain the initial values for the Bethe roots by computing the eigenvalues of the transfer matrix $T^{(c)}(x)$ as meromorphic functions on $x$. We use this data to solve the Bethe equation (15) with high numerical precision and the energy and the spin of the states are calculated with the help of Eqs.(16,51). The energies obtained from the Bethe roots are then checked with those coming from the exact diagonalization of the Hamiltonian $H^{(c)}$. The results are shown in table (2).
$E^{(c)}$ $\nu_{c}$ $p$ $\lambda_{j}$
-5.77350269 +1 0 $0.13588376\pm 0.57005210i$$0.13588376\pm 0.57005210i$
-3.89397505 +1 0 $\pm 0.50828016i,~\pm 0.524244017$
[…]
1.154700538 +1 11 $0.28844398\pm 0.61762004i$0.558766300.558766300.018121660.01812166
1.154700538 +1 1-1 $0.28844398\pm 0.61762004i$0.55876630-0.558766300.01812166-0.01812166
2.739274521 +1 0 $\pm 1.08877871i$$\pm 0.14111440$
[…]
Table 2: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=2$. For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$).
@@ 5 Other families of integrable Hamiltonians
$\mathrm{\tilde{T}}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}1}(x)\right],$ (53)

### v2

@@ 4 The Bethe equation for $T_{ver}^{(c)}(x)$
where now the number of sine terms is the same as that present in the general form (34) for the transfer matrix elements. We arrive at this conclusion by inspecting the asymptotic limit of the eigenvalues of $T_{ver}^{(c)}(x)$ at the points $\pm i\infty$ for small lattice sizes.
As before, the Bethe equation for the variables $\xi_{k}$ is uncovered with the help of a functional relation for the eigenvalue involving distinct spectral parameters. This equation is derived from a matrix identity for the product of transfer matrices whose main structure is similar to that presented in previous section, see Eq.(38). More specifically, we find that $T_{ver}^{(c)}(x)$ satisfies the following matrix identity,
$T_{ver}^{(c)}(x-\frac{\pi}{3})T_{ver}^{(c)}(x-\frac{\pi}{6})T_{ver}^{(c)}(x)$ $=$ $T_{ver}^{(c)}(0)\bigg\{\big[f_{1}(x)\big]^{L}T_{ver}^{(c)}(x-\frac{\pi}{3})+\big[f_{2}(x)\big]^{L}T_{ver}^{(c)}(x)$ (49)
- $\big[f_{3}(x)\big]^{L}T{ver}^{(c)}(x+\frac{\pi}{3})\bigg\}$
where the main difference with Eq.(38) is the presence of a minus sign on the term proportional to the amplitude $f_{3}(x)$. We have established this relation using the same procedure explained in the previous section. We have checked that such matrix relation is satisfied up to $L=5$ and we conjecture that it is valid for arbitrary $L$.
@@ 4.1 Bethe roots for L=2
We now proceed to investigate the Bethe equation parameterization we have found for the spectrum of the Hamiltonian $H^{(c)}$ with $L=2$. This Hamiltonian has a $Z(2)$ invariance and the Hilbert space splits into two sectors labeled by the eigenvalues of the charge ${\cal{O}}_{Z(2)}$. If we denote these eigenvalues by the symbol $\nu_{c}$ we find that there are $(3^{L}+1)/2$ states with $\nu_{c}=+1$ and $(3^{L}-1)/2$ states with $\nu_{c}=-1$. We obtain the initial values for the Bethe roots by computing the eigenvalues of the transfer matrix $T_{ver}^{(c)}(x)$ as meromorphic functions on $x$. We use this data to solve the Bethe equation (15) with high numerical precision and the energy and the spin of the states are calculated with the help of Eqs.(16,52). The energies obtained from the Bethe roots are then checked with those coming from the exact diagonalization of the Hamiltonian $H^{(c)}$. The results are shown in table (2). In this table we have also presented the data for the corresponding transfer matrix eigenvalue $\Lambda^{(c)}(x)$ at $x=\frac{3}{5}$ computed by using Eqs.(43,48). We stress that these transfer matrix eigenvalues agree with the values obtained from the diagonalization of the operator $T_{ver}^{(c)}(\frac{3}{5})$.
$E^{(c)}$ $\nu_{c}$ $p$ $\lambda_{j}$ $\Lambda^{(c)}(\frac{3}{5})$
-5.77350269 +1 0 $\begin{array}[c]{c}0.13588376\pm 0.57005210i,\\ -0.13588376\pm 0.57005210i\end{array}$ $0.90168671$
-3.89397505 +1 0 $\pm 0.50828016i,~\pm 0.524244017$ $1.02818296$
-1.67372658 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}0.13542121\pm 0.5300866i,\\ -0.14584486,-0.12499756+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.23423027\\ -0.06662064i\end{array}$
-1.67372658 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}-0.13542121\pm 0.5300866i,\\ 0.14584486,~0.12499756+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.23423027\\ +0.06662064i\end{array}$
1.154700538 +1 11 $\begin{array}[c]{c}-0.28844398\pm 0.61762004i,\\ 0.55876630,~0.01812166\end{array}$ $\begin{array}[c]{c}1.54404018\\ +0.03665078i\end{array}$
1.154700538 +1 1-1 $\begin{array}[c]{c}0.28844398\pm 0.61762004i,\\ -0.55876630,~-0.01812166\end{array}$ $\begin{array}[c]{c}1.54404018\\ -0.03665078i\end{array}$
2.739274521 +1 0 $\pm 1.08877871i$$\pm 0.14111440$ $1.73331161$
3.983127663 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}-0.59441775,~-0.05108087,\\ ~0.19398856,~0.45151005+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.91132869\\ +0.00914203i\end{array}$
3.983127663 1-1 $\frac{1}{2}$ $\begin{array}[c]{c}0.59441775$,~$0.05108087,\\ -0.19398856,~-0.45151005+i\frac{\pi}{2}\end{array}$ $\begin{array}[c]{c}1.91132869\\ -0.00914203i\end{array}$
Table 2: The eigenvalues ($E^{(c)}$) of the Hamiltonian $H^{(c)}$ (14) for $L=2$. For each energy, we indicate the sector ($\nu_{c}$), the spin of the state $(s_{p})$ and the values of the Bethe roots ($\lambda_{j}$). We have also added the respective transfer matrix eigenvalue $\Lambda^{(c)}(x)$ at $x=\frac{3}{5}$.
@@ 5 Other families of integrable Hamiltonians
$\tilde{T}_{ver}(x)=\mathrm{Tr}_{\cal{A}}\left[\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L}(x)\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}L-1}(x)\dots\mathrm{G}_{\cal{A}}\mathbb{L}_{{\cal{A}}1}(x)\right],$ (54)